OpenStudy (anonymous):

Solve -4(-2y + 3) = 20. -1 1 4 -4

OpenStudy (anonymous):

1

OpenStudy (anonymous):

multiply both sides by -4 then subtract 3 on both sides then divide by2 in both sides

OpenStudy (amistre64):

divide .... not multiply

OpenStudy (amistre64):

and its a -2

OpenStudy (amistre64):

other than that, spot on :)

mathslover (mathslover):

\[\large{-4(-2y+3)=20}\] \[\large{\frac{-4(-2y+3)}{-4}=\frac{20}{-4}}\]

OpenStudy (jiteshmeghwal9):

Distribute -4 btw -2y+3 -4(-2y)+(-4(3))=20 8y-12=20 8y=20+12 y=32/8=?????

OpenStudy (amistre64):

distributing the -4 i spose is a way to go, but pointless since its not really attached to a y to begin with

OpenStudy (anonymous):

4

OpenStudy (jiteshmeghwal9):

yes! @miadeex